# FragmentGPT: A Unified GPT Model for Fragment Growing, Linking, and Merging in Molecular Design

### v1

@@ 1 Introduction
Fragment-based drug design (FBDD) has become a mainstream strategy in early drug discovery, especially for challenging targets like protein–protein interactions, allosteric sites, and shallow or cryptic pockets (Xu and Kang, 2025). Unlike high-throughput screening (HTS), which depends on large, complex compound libraries (Leach and Hann, 2011), FBDD employs small, low-molecular-weight fragments to probe binding surfaces with precision and efficiency (Keserű et al., 2016; Leach and Hann, 2011; Xu and Kang, 2025; Murray and Rees, 2009). Though fragments typically bind weakly (micromolar–millimolar range), they offer high ligand efficiency and chemical tractability, making them ideal starting points for optimization (Wasko et al., 2015). In addition, fragment libraries, usually only hundreds to a few thousand compounds, cover chemical space more effectively than HTS due to the combinatorial potential of linking or growing fragments. Their small, hydrophilic nature often yields hits with favorable solubility and reduced off-target risks. Moreover, sensitive biophysical methods such as NMR, X-ray crystallography, SPR, and thermal shift assays enable rapid screening and validation, making FBDD a powerful, efficient route for lead discovery.
The impact of FBDD is underscored by the approval of several clinically significant drugs. For example, Zelboraf (vemurafenib), a BRAF inhibitor for melanoma, was the first FDA-approved drug derived from FBDD (Bollag et al., 2012). It was followed by others such as Venetoclax, a Bcl-2 inhibitor disrupting protein–protein interactions (Fairbrother et al., 2019), and Sotorasib, a KRAS-G12C inhibitor targeting a previously “undruggable” oncogene (Lanman et al., 2020). In total, eight FDA-approved drugs and over 50 clinical candidates have originated from FBDD pipelines, validating its versatility across diverse target classes, including kinases, protein–protein interactions, and even RNA-binding proteins (Xu and Kang, 2025).
@@
Fragment growing involves expanding a single fragment into nearby binding regions, guided by structural information or structure–activity relationships (SAR). The clinical success of fragment growing is exemplified by several notable drug discoveries. Smith and colleagues advanced MRTX1719, targeting the PRMT5·MTA complex, beginning with 4-(aminomethyl)quinazoline-1(2H) identified through surface plasmon resonance (Smith et al., 2022). Similarly, Heinrich and colleagues developed focal adhesion kinase inhibitors starting from 1H-pyrrolo[2,3-b]pyridine scaffolds (Heinrich et al., 2013), while Saxty et al. utilized their "Pyramid" X-ray crystallography platform to evolve 5-methyl-4-phenyl-1H-pyrazole hits into potent protein kinase B inhibitors (Wyatt et al., 2008). Likewise, Addie and colleagues at AstraZeneca started with a small hinge-binding pyrrolopyrimidine fragment and, using crystal structures to guide each tweak, built it into AZD5363 (capivasertib)—an oral, ATP-competitive Akt inhibitor with broad isoform potency, better ROCK selectivity, low hERG liability, and strong anti-tumor activity in cells and animals (Addie et al., 2013). Fragment growing offers distinct advantages including efficient exploration of chemical space through incremental expansion, structure-guided optimization based on detailed binding insights, and typically higher ligand efficiency compared to traditional high-throughput screening approaches. The method enables rational design decisions by leveraging crystallographic or NMR structural information to direct growth toward productive binding site regions.
@@
Fragment merging integrates two or more overlapping fragments with shared substructures or binding features to preserve favorable interactions and enhance potency. Traditionally, this requires crystal structures to pinpoint overlaps—a labor-intensive process demanding expert knowledge—and even then, maintaining bioactive conformations often necessitates extensive optimization. Successful examples include trypanothione reductase inhibitors, where a propylphenyl and p-fluorophenyl group were merged via a piperazine scaffold (Exertier et al., 2024), and HSP90 inhibitors, where adjacent fragments sharing a phenol moiety were combined but initially showed weak activity due to desolvation penalties (Ren et al., 2014). Computational tools like the Fragment Network extend this idea by searching over 120 million catalogue compounds to suggest merges beyond simple fingerprint similarity (Wills et al., 2023), offering access to new regions of chemical space. However, such methods remain constrained by database coverage and lack true generative flexibility, leaving fragment merging an appealing yet underdeveloped strategy in practice.
@@
Recent advances in machine learning have introduced generative models as powerful tools for automated linker design. Sequence-based models, such as SyntaLinker (Yang et al., 2020), formulate linker generation as a sequence-to-sequence task, treating linkers as SMILES strings and employing natural language processing techniques, particularly Transformer-based neural networks. These models offer fast inference and strong chemical priors, effectively capturing the syntactic and semantic regularities of chemical representations. However, their reliance on 1D chemical language often necessitates post-processing steps to ensure the 3D structural feasibility and compatibility of generated linkers with protein binding pockets. In contrast, 3D geometry-based models, including those based on diffusion processes and variational autoencoders (VAEs)—such as 3DLinker (Huang et al., 2022), LinkerNet (Guan et al., 2023), DiffLinker (Igashov et al., 2024), and DiffPROTACs (Li et al., 2024) —generate linkers directly in three-dimensional space. These models exploit E(3)-equivariant architectures that inherently preserve rotational and translational symmetries, enabling the generation of spatially compatible and structurally plausible linkers conditioned on the geometric context of fragment binding poses and protein pockets.
@@
In particular, DiffLinker, a state-of-the-art E(3)-equivariant 3D-conditional diffusion model, has demonstrated impressive performance in generating chemically valid, diverse, and pocket-compatible linkers between arbitrary numbers of fragments (Igashov et al., 2024). By modeling atomic geometries directly and conditioning on both fragment coordinates and protein pockets, DiffLinker surpasses prior approaches based on SMILES or autoregressive graph construction in terms of synthetic accessibility, drug-likeness, and similarity of samples to reference compounds. However, despite these achievements, current methods—including DiffLinker—focus solely on linking distinct fragments and remain limited in handling fragment merging, where two overlapping or partially redundant fragments must be merged into a single coherent substructure. No existing generative framework can yet simultaneously address both the linking and merging tasks in fragment-based design. Developing such unified strategies is critical for enabling more flexible and realistic drug design workflows where overlapping binding poses or scaffold overlaps frequently occur.
@@
Currently, no generative framework can jointly address both fragment linking and merging and growing. The rise of large language models (LLMs) introduces a powerful paradigm for generative chemistry, yet their potential in fragment-based design—particularly linking and merging—remains largely underexplored. Moreover, pretrained GPT models rely solely on maximum likelihood training, lacking explicit objective-driven optimization. What’s more, pretrained GPT models have shown encouraging results in molecular tasks, highlighting the opportunity to harness these expert models (Liu et al., 2022, 2023b) to boost the performance of our framework. Finally, biomolecular design is inherently a multi-objective optimization problem. Multi-property evaluation measures how well a model balances multiple, often conflicting, objectives at once. One approach is a composite score, where different property metrics (e.g., LogP, QED, synthetic accessibility) are combined into a single weighted objective, guiding optimization methods. Another is Pareto optimality, where solutions are considered optimal if no property can be improved without worsening another; the collection of these solutions forms the Pareto front, which helps visualize trade-offs among objectives. However, both approaches come with their own advantages and limitations. To address these challenges, we asked
@@ 2 Preliminaries
LM. Let ${F}=\left[{f}_{1},{f}_{2},\cdots,{f}_{n}\right]$ denote an input sequence of tokens (fragment-based prompt), where each ${f}_{i}$ belongs to a vocabulary $\mathcal{V}$. The output sequence is defined as $M=\left[m_{1},m_{2},\cdots,m_{T}\right]$ with $m_{i}\in\mathcal{M}$, where $\mathcal{V}$ and $\mathcal{M}$ may be distinct vocabularies. For convenience, we write $\mathbf{m}{<t}=\left[m_{1},\cdots,m{t-1}\right]$ and $\mathbf{m}_{T}:=M$, and use $T$ to denote sequence length. Each training sample begins with a start token $\left[\text{BOS}\right]$, followed by a sequence $\mathbf{m}$ with $y_{i}\in\mathcal{V}$, and terminates with an end token $\left[\text{EOS}\right]$. Molecules are represented as token sequences $\mathbf{m}$ that assemble into SMILES strings, covering both partial and complete structures. We denote string concatenation by $\circ$ and the Kleene closure of $\mathcal{V}$ by $\mathcal{V}^{*}$. The training corpus $C$ is thus: $C:=\left\{\text{[BOS]}\circ\mathbf{v}\circ\text{[EOS]}~|~\mathbf{v}\in\mathcal{V}^{*}\right\}.$ The generator policy $\pi_{\theta}$, parameterized by a deep neural network with weights $\theta$, defines a distribution over sequences as $\pi_{\theta}\left({\mathbf{m}|\mathbf{f}}\right)=\prod_{t=1}^{|\mathbf{m}|}\pi_{\theta}\left({m_{t}|\mathbf{f},\mathbf{m}_{<t}}\right)$, where $\pi_{\theta}(m_{t}|\mathbf{f},\mathbf{m}{<t})=P(m_{t}|\mathbf{m}{<t},X)$ denotes the conditional probability of token $m_{t}$. Decoding aims to select the most likely sequence from the hypothesis space: $\mathbf{m}^{\star}=\mathop{\rm arg\,max}_{\mathbf{m}\in\mathcal{Y}_{T}}\log\pi_{\theta}\left({\mathbf{m}|\mathbf{f}}\right).$ Causal language modeling (CLM) reformulates the training objective as $\max_{\theta}\sum^{n}_{i=1}\log P\left({f_{i}|{{F}}_{<i};\theta}\right)$, where $P\left({f_{i}|{{F}}_{<i};\theta}\right)$, where $P(f_{i}|{F}_{<i};\theta)$ is the conditional probability of token ${f}_{i}$ given its history. In this work, we employ a GPT policy $\pi_{\theta}$ trained on dataset $\mathcal{B}$ to capture prior chemical knowledge and generate chemically valid molecules. Molecules are represented using SMILES strings (Weininger, 1988). The tokenizer is first pre-trained with the Byte Pair Encoding (BPE) algorithm (Gage, 1994; Sennrich et al., 2015). Building on this, we adopt an incremental pre-training strategy inspired by ScaffoldGPT (Liu et al., 2025b), which significantly enhances the model’s ability to reconstruct valid molecules from fragments.
@@ 3 The FragmentGPT Algorithm
Here, $T$ denotes the total number of tokens in the target molecule $Y$. The negative log-likelihood (NLL) thus measures the likelihood of reconstructing a designated target molecule from the given fragment pair $\left\{F_{i=1}^{n}\right\}$. In this project, we adopt an incremental training strategy (Liu et al., 2025b) that leverages large collections of unlabeled text to build a foundational model of language understanding. This foundation model can then be fine-tuned and adapted to a variety of downstream tasks with specific objectives.
@@
Stage 2. Expert Exploration. We begin by sampling a batch of prompts $\mathcal{X}_{t}=\left\{{\left\{F_{i=1}^{n}\right\}}_{1}^{t},\cdots,{\left\{F_{i=1}^{n}\right\}}_{b}^{t}\right\}\subseteq\mathcal{X}$, and for each ${\left\{F_{i=1}^{n}\right\}}_{i}^{t}\in\mathcal{X}_{t}$, we generate molecules $M_{1},\cdots,M_{K}\sim\pi_{\theta}\left({\cdot|{\left\{F_{i=1}^{n}\right\}}_{i}^{t}}\right)$. Next, we leverage the expert policy $\pi^{\textsc{E}}$ to produce more diverse molecular candidates: $M_{1}^{\textsc{Expert}},\cdots,M_{K}^{\textsc{Expert}}\sim\pi^{\textsc{Expert}}\left({\cdot|M_{i}}\right),i\in\left[K\right]$. Specifically, we employ ScaffoldGPT (Liu et al., a, 2025b) and ControllableGPT (Liu et al., 2025c, b) to promote structural diversity. Finally, we combine the molecules generated by both the learner policy and the expert policy into the replay buffer $\mathcal{B}$, $\mathcal{B}\leftarrow\left\{M\right\}\cup\left\{M^{\textsc{Expert}}\right\}$.
Stage 3. Data Selection and Augmentation. In this step, we employ the Pareto optimality reward model $R^{\textsc{Pareto}}\left({M}\right)$} from Abeer et al. (2024) and the composite score reward model $R^{\textsc{Composite}}\left({M}\right)$} from Liu et al. (2025b). In this work, we regarded each pharmaceutical property as a critic $C$ and considered an ensemble critics $\mathbf{C}$ as follows:
@@
by randomly sampling 20% from the previous $\mathcal{D}$ and 80% from the buffer, with sampling prioritized according to their ranking (Liu et al., 2025d). We then proceed to Stage 1, where we retrain the learner policy on the dataset aligned with Pareto optimality and composite score.
@@ 4.1 Experimental configuration
Baselines. We benchmark our approach against baseline model, Link-INVENT (Guo et al., 2023), which employs an encoder–decoder architecture based on recurrent neural networks (RNNs). In this framework, two molecular fragments, referred to as warheads, are provided as input constraints to the RNN-based generator, which then produces a suitable linker to connect them.
Dataset. We utilize 1 million compounds from the ZINC15 database, docked against the human cancer protein RTCB (PDB ID: 4DWQ), as provided in the most recent Cancer dataset (Liu et al., 2023a, 2025a).
Critics and evaluation metric. In this work, we assess the effectiveness of FragmentGPT in generating molecules that exhibit favorable characteristics for pharmaceutical drug discovery. To perform this evaluation, we employ the RDKit chemoinformatics toolkit (Landrum et al., ) and consider the following performance metrics. Validity determines whether the generated SMILES strings are syntactically correct. Druglikeness evaluates the probability that a molecule could serve as a viable drug candidate. Solubility estimates the likelihood that a molecule can dissolve in water, approximated by the water–octanol partition coefficient (LogP). Synthesizability measures the practical feasibility of chemical synthesis, where a score of 1 indicates high ease of synthesis and 10 indicates high difficulty (Ertl and Schuffenhauer, 2009). Docking Score reflects the binding potential of a molecule to the target site; for computational efficiency, this score is predicted using a surrogate docking model (see Appendix A.3). Similarity is quantified by the Tanimoto coefficient, comparing the structural similarity between the generated SMILES and the original molecules. Average Top 10% Normalized Reward captures the mean normalized reward across the top 10% of molecules, ranked by their normalized reward. Average Normalized Reward, our primary evaluation metric, represents the mean normalized values of docking score, druglikeness, synthesizability, solubility, and similarity over all valid molecules. All evaluations are conducted on a sample of 1,280 molecules.

### v2

@@ 1 Introduction
Fragment-based drug design (FBDD) has become a mainstream strategy in early drug discovery, especially for challenging targets like protein–protein interactions, allosteric sites, and shallow or cryptic pockets [39]. Unlike high-throughput screening (HTS), which depends on large, complex compound libraries [18], FBDD employs small, low-molecular-weight fragments to probe binding surfaces with precision and efficiency [15, 18, 39, 29]. Though fragments typically bind weakly (micromolar–millimolar range), they offer high ligand efficiency and chemical tractability, making them ideal starting points for optimization [34]. In addition, fragment libraries, usually only hundreds to a few thousand compounds, cover chemical space more effectively than HTS due to the combinatorial potential of linking or growing fragments. Their small, hydrophilic nature often yields hits with favorable solubility and reduced off-target risks. Moreover, sensitive biophysical methods such as NMR, X-ray crystallography, SPR, and thermal shift assays enable rapid screening and validation, making FBDD a powerful, efficient route for lead discovery.
The impact of FBDD is underscored by the approval of several clinically significant drugs. For example, Zelboraf (vemurafenib), a BRAF inhibitor for melanoma, was the first FDA-approved drug derived from FBDD [4]. It was followed by others such as Venetoclax, a Bcl-2 inhibitor disrupting protein–protein interactions [7], and Sotorasib, a KRAS-G12C inhibitor targeting a previously “undruggable” oncogene [17]. In total, eight FDA-approved drugs and over 50 clinical candidates have originated from FBDD pipelines, validating its versatility across diverse target classes, including kinases, protein–protein interactions, and even RNA-binding proteins [39].
@@
Fragment growing involves expanding a single fragment into nearby binding regions, guided by structural information or structure–activity relationships (SAR). The clinical success of fragment growing is exemplified by several notable drug discoveries. Smith and colleagues advanced MRTX1719, targeting the PRMT5·MTA complex, beginning with 4-(aminomethyl)quinazoline-1(2H) identified through surface plasmon resonance [32]. Similarly, Heinrich and colleagues developed focal adhesion kinase inhibitors starting from 1H-pyrrolo[2,3-b]pyridine scaffolds [12], while Saxty et al. utilized their "Pyramid" X-ray crystallography platform to evolve 5-methyl-4-phenyl-1H-pyrazole hits into potent protein kinase B inhibitors [38]. Likewise, Addie and colleagues at AstraZeneca started with a small hinge-binding pyrrolopyrimidine fragment and, using crystal structures to guide each tweak, built it into AZD5363 (capivasertib)—an oral, ATP-competitive Akt inhibitor with broad isoform potency, better ROCK selectivity, low hERG liability, and strong anti-tumor activity in cells and animals [2]. Fragment growing offers distinct advantages including efficient exploration of chemical space through incremental expansion, structure-guided optimization based on detailed binding insights, and typically higher ligand efficiency compared to traditional high-throughput screening approaches. The method enables rational design decisions by leveraging crystallographic or NMR structural information to direct growth toward productive binding site regions.
@@
Fragment merging integrates two or more overlapping fragments with shared substructures or binding features to preserve favorable interactions and enhance potency. Traditionally, this requires crystal structures to pinpoint overlaps—a labor-intensive process demanding expert knowledge—and even then, maintaining bioactive conformations often necessitates extensive optimization. Successful examples include trypanothione reductase inhibitors, where a propylphenyl and p-fluorophenyl group were merged via a piperazine scaffold [6], and HSP90 inhibitors, where adjacent fragments sharing a phenol moiety were combined but initially showed weak activity due to desolvation penalties [30]. Computational tools like the Fragment Network extend this idea by searching over 120 million catalogue compounds to suggest merges beyond simple fingerprint similarity [36], offering access to new regions of chemical space. However, such methods remain constrained by database coverage and lack true generative flexibility, leaving fragment merging an appealing yet underdeveloped strategy in practice.
@@
Recent advances in machine learning have introduced generative models as powerful tools for automated linker design. Sequence-based models, such as SyntaLinker [41], formulate linker generation as a sequence-to-sequence task, treating linkers as SMILES strings and employing natural language processing techniques, particularly Transformer-based neural networks. These models offer fast inference and strong chemical priors, effectively capturing the syntactic and semantic regularities of chemical representations. However, their reliance on 1D chemical language often necessitates post-processing steps to ensure the 3D structural feasibility and compatibility of generated linkers with protein binding pockets. In contrast, 3D geometry-based models, including those based on diffusion processes and variational autoencoders (VAEs)—such as 3DLinker [13], LinkerNet [10], DiffLinker [14], and DiffPROTACs [19] —generate linkers directly in three-dimensional space. These models exploit E(3)-equivariant architectures that inherently preserve rotational and translational symmetries, enabling the generation of spatially compatible and structurally plausible linkers conditioned on the geometric context of fragment binding poses and protein pockets.
@@
In particular, DiffLinker, a state-of-the-art E(3)-equivariant 3D-conditional diffusion model, has demonstrated impressive performance in generating chemically valid, diverse, and pocket-compatible linkers between arbitrary numbers of fragments [14]. By modeling atomic geometries directly and conditioning on both fragment coordinates and protein pockets, DiffLinker surpasses prior approaches based on SMILES or autoregressive graph construction in terms of synthetic accessibility, drug-likeness, and similarity of samples to reference compounds. However, despite these achievements, current methods—including DiffLinker—focus solely on linking distinct fragments and remain limited in handling fragment merging, where two overlapping or partially redundant fragments must be merged into a single coherent substructure. No existing generative framework can yet simultaneously address both the linking and merging tasks in fragment-based design. Developing such unified strategies is critical for enabling more flexible and realistic drug design workflows where overlapping binding poses or scaffold overlaps frequently occur.
@@
Currently, no generative framework can jointly address both fragment linking and merging and growing. The rise of large language models (LLMs) introduces a powerful paradigm for generative chemistry, yet their potential in fragment-based design—particularly linking and merging—remains largely underexplored. Moreover, pretrained GPT models rely solely on maximum likelihood training, lacking explicit objective-driven optimization. What’s more, pretrained GPT models have shown encouraging results in molecular tasks, highlighting the opportunity to harness these expert models [22, 24] to boost the performance of our framework. Finally, biomolecular design is inherently a multi-objective optimization problem. Multi-property evaluation measures how well a model balances multiple, often conflicting, objectives at once. One approach is a composite score, where different property metrics (e.g., LogP, QED, synthetic accessibility) are combined into a single weighted objective, guiding optimization methods. Another is Pareto optimality, where solutions are considered optimal if no property can be improved without worsening another; the collection of these solutions forms the Pareto front, which helps visualize trade-offs among objectives. However, both approaches come with their own advantages and limitations. To address these challenges, we asked
@@ 2 Preliminaries
LM. Let ${F}=\left[{f}_{1},{f}_{2},\cdots,{f}_{n}\right]$ denote an input sequence of tokens (fragment-based prompt), where each ${f}_{i}$ belongs to a vocabulary $\mathcal{V}$. The output sequence is defined as $M=\left[m_{1},m_{2},\cdots,m_{T}\right]$ with $m_{i}\in\mathcal{M}$, where $\mathcal{V}$ and $\mathcal{M}$ may be distinct vocabularies. For convenience, we write $\mathbf{m}{<t}=\left[m_{1},\cdots,m{t-1}\right]$ and $\mathbf{m}_{T}:=M$, and use $T$ to denote sequence length. Each training sample begins with a start token $\left[\text{BOS}\right]$, followed by a sequence $\mathbf{m}$ with $y_{i}\in\mathcal{V}$, and terminates with an end token $\left[\text{EOS}\right]$. Molecules are represented as token sequences $\mathbf{m}$ that assemble into SMILES strings, covering both partial and complete structures. We denote string concatenation by $\circ$ and the Kleene closure of $\mathcal{V}$ by $\mathcal{V}^{*}$. The training corpus $C$ is thus: $C:=\left\{\text{[BOS]}\circ\mathbf{v}\circ\text{[EOS]}~|~\mathbf{v}\in\mathcal{V}^{*}\right\}.$ The generator policy $\pi_{\theta}$, parameterized by a deep neural network with weights $\theta$, defines a distribution over sequences as $\pi_{\theta}\left({\mathbf{m}|\mathbf{f}}\right)=\prod_{t=1}^{|\mathbf{m}|}\pi_{\theta}\left({m_{t}|\mathbf{f},\mathbf{m}_{<t}}\right)$, where $\pi_{\theta}(m_{t}|\mathbf{f},\mathbf{m}{<t})=P(m_{t}|\mathbf{m}{<t},X)$ denotes the conditional probability of token $m_{t}$. Decoding aims to select the most likely sequence from the hypothesis space: $\mathbf{m}^{\star}=\mathop{\rm arg\,max}_{\mathbf{m}\in\mathcal{Y}_{T}}\log\pi_{\theta}\left({\mathbf{m}|\mathbf{f}}\right).$ Causal language modeling (CLM) reformulates the training objective as $\max_{\theta}\sum^{n}_{i=1}\log P\left({f_{i}|{{F}}_{<i};\theta}\right)$, where $P\left({f_{i}|{{F}}_{<i};\theta}\right)$, where $P(f_{i}|{F}_{<i};\theta)$ is the conditional probability of token ${f}_{i}$ given its history. In this work, we employ a GPT policy $\pi_{\theta}$ trained on dataset $\mathcal{B}$ to capture prior chemical knowledge and generate chemically valid molecules. Molecules are represented using SMILES strings [35]. The tokenizer is first pre-trained with the Byte Pair Encoding (BPE) algorithm [9, 31]. Building on this, we adopt an incremental pre-training strategy inspired by ScaffoldGPT [26], which significantly enhances the model’s ability to reconstruct valid molecules from fragments.
@@ 3 The FragmentGPT Algorithm
Here, $T$ denotes the total number of tokens in the target molecule $Y$. The negative log-likelihood (NLL) thus measures the likelihood of reconstructing a designated target molecule from the given fragment pair $\left\{F_{i=1}^{n}\right\}$. In this project, we adopt an incremental training strategy [26] that leverages large collections of unlabeled text to build a foundational model of language understanding. This foundation model can then be fine-tuned and adapted to a variety of downstream tasks with specific objectives.
@@
Stage 2. Expert Exploration. We begin by sampling a batch of prompts $\mathcal{X}_{t}=\left\{{\left\{F_{i=1}^{n}\right\}}_{1}^{t},\cdots,{\left\{F_{i=1}^{n}\right\}}_{b}^{t}\right\}\subseteq\mathcal{X}$, and for each ${\left\{F_{i=1}^{n}\right\}}_{i}^{t}\in\mathcal{X}_{t}$, we generate molecules $M_{1},\cdots,M_{K}\sim\pi_{\theta}\left({\cdot|{\left\{F_{i=1}^{n}\right\}}_{i}^{t}}\right)$. Next, we leverage the expert policy $\pi^{\textsc{E}}$ to produce more diverse molecular candidates: $M_{1}^{\textsc{Expert}},\cdots,M_{K}^{\textsc{Expert}}\sim\pi^{\textsc{Expert}}\left({\cdot|M_{i}}\right),i\in\left[K\right]$. Specifically, we employ ScaffoldGPT [20, 26] and ControllableGPT [27, 21] to promote structural diversity. Finally, we combine the molecules generated by both the learner policy and the expert policy into the replay buffer $\mathcal{B}$, $\mathcal{B}\leftarrow\left\{M\right\}\cup\left\{M^{\textsc{Expert}}\right\}$.
Stage 3. Data Selection and Augmentation. In this step, we employ the Pareto optimality reward model $R^{\textsc{Pareto}}\left({M}\right)$} from Abeer et al. [1] and the composite score reward model $R^{\textsc{Composite}}\left({M}\right)$} from Liu et al. [26]. In this work, we regarded each pharmaceutical property as a critic $C$ and considered an ensemble critics $\mathbf{C}$ as follows:
@@
by randomly sampling 20% from the previous $\mathcal{D}$ and 80% from the buffer, with sampling prioritized according to their ranking [28]. We then proceed to Stage 1, where we retrain the learner policy on the dataset aligned with Pareto optimality and composite score.
@@ 4.1 Experimental configuration
Baselines. We benchmark our approach against baseline model, Link-INVENT [11], which employs an encoder–decoder architecture based on recurrent neural networks (RNNs). In this framework, two molecular fragments, referred to as warheads, are provided as input constraints to the RNN-based generator, which then produces a suitable linker to connect them.
Dataset. We utilize 1 million compounds from the ZINC15 database, docked against the human cancer protein RTCB (PDB ID: 4DWQ), as provided in the most recent Cancer dataset [23, 25].
Critics and evaluation metric. In this work, we assess the effectiveness of FragmentGPT in generating molecules that exhibit favorable characteristics for pharmaceutical drug discovery. To perform this evaluation, we employ the RDKit chemoinformatics toolkit [16] and consider the following performance metrics. Validity determines whether the generated SMILES strings are syntactically correct. Druglikeness evaluates the probability that a molecule could serve as a viable drug candidate. Solubility estimates the likelihood that a molecule can dissolve in water, approximated by the water–octanol partition coefficient (LogP). Synthesizability measures the practical feasibility of chemical synthesis, where a score of 1 indicates high ease of synthesis and 10 indicates high difficulty [5]. Docking Score reflects the binding potential of a molecule to the target site; for computational efficiency, this score is predicted using a surrogate docking model (see Appendix A.3). Similarity is quantified by the Tanimoto coefficient, comparing the structural similarity between the generated SMILES and the original molecules. Average Top 10% Normalized Reward captures the mean normalized reward across the top 10% of molecules, ranked by their normalized reward. Average Normalized Reward, our primary evaluation metric, represents the mean normalized values of docking score, druglikeness, synthesizability, solubility, and similarity over all valid molecules. All evaluations are conducted on a sample of 1,280 molecules.